# Memorization Over Reasoning? Exposing and Mitigating Verbatim Memorization in Large Language Models’ Character Understanding Evaluation

### v3

@@ Abstract
Recently, Large Language Models (LLMs) have shown impressive performance in character understanding tasks, such as analyzing the roles, personalities, and relationships of fictional characters. However, the extensive pre-training corpora used by LLMs raise concerns that they may rely on memorizing popular fictional works rather than genuinely understanding and reasoning about them. In this work, we argue that “gist memory”—capturing essential meaning—should be the primary mechanism for character understanding tasks, as opposed to “verbatim memory”—exact match of a string. We introduce a simple yet effective method to mitigate mechanized memorization in character understanding evaluations while preserving the essential implicit cues needed for comprehension and reasoning. Our approach reduces memorization-driven performance on popular fictional works from 96% accuracy to 72% and results in up to an 18% drop in accuracy across various character understanding tasks. These findings underscore the issue of data contamination in existing benchmarks, which often measure memorization rather than true character understanding.
@@ 1 Introduction
Motivated by these concepts, we propose an approach that alters elements such as character names, place names, and specific notations in the sentences while preserving the relationships between characters, significant events, and other information we want the model to utilize. We hypothesize that these modifications can reduce the model’s tendency to take shortcuts by relying on verbatim memory, and instead encourages the model to rely on gist memory to stimulate its reasoning ability for character understanding tasks. We tested our method on six key types of character understanding tasks and obtained an interesting finding: When the difficulty of using verbatim memory increases, the model is more likely to resort to reasoning to compensate.
Our work provides a simple approach to ensure a fairer benchmark for character understanding tasks on large models that already have strong memorization abilities of fictional works. It offers an alternative method compared to more costly solutions, such as developing data from obscure or very new books that the model has not encountered Ngo et al. (2024), or using techniques like machine unlearning to forget original training texts Shaik et al. (2024).
In summary, we provide an outline of the paper and summarize our key contributions:
  • To determine whether the memory types used in character understanding tasks align with our expectations, we draw on human cognitive science to define gist and verbatim memory in the context of character understanding tasks (s 2).
  • We benchmarked the mechanism memorization ability of over 100 sources of fictional works (s 4.1).
  • We measured the verbatim memory of large models for hundreds of different source fiction works and tested how different keywords affect the verbatim memory of these models (s 4).
  • We demonstrated that even for the same factor, such as a person’s name, different disturbances, such as gender or cultural interference, can lead to significantly different verbatim memory effects in large models (s 5).
  • We compared the different effects of the model using verbatim and gist memory across six different test sets and found that the model’s accuracy decreased by 0.7% to 17.2% (s 6).
2 Related Task
Among Large Language Models (LLMs)’ various capabilities Li et al. (2024b, d); Zhang et al. (2024); Li et al. (2024c), the reasoning abilities Tong et al. (2023); Yang et al. (2023); Xiong et al. (2024b); Tong et al. (2024); Xiong et al. (2024a) of LLMs Li et al. (2024a); Tan et al. (2024) have garnered significant attention, but their memory has also been shown to play a crucial role in influencing benchmark reasoning performance Zhou et al. (2023). This influence is complex: data contamination can inflate performance metrics Khandelwal et al. (2020); Blevins and Zettlemoyer (2022); Carlini et al. (2023), while knowledge conflict can reduce them Su et al. (2024); Wang et al. (2024); Xie et al. (2024b); Xu et al. (2024). Recent efforts have aimed to mitigate the impact of memory on reasoning benchmarks to establish more reliable evaluations, particularly for reasoning tasks like numerical reasoning, commonsense reasoning, or logical reasoning, where memory requirements are less significant and easier to define Xie et al. (2024a); Sainz et al. (2023); Li et al. (2024e). However, in character understanding tasks, the diversity of task formats and scenarios means that memory requirements vary significantly, making it challenging to establish a universal standard to distinguish between acceptable and unacceptable use of memory. We propose leveraging gist memory to reduce LLMs’ reliance on precise memorization. This approach encourages models to utilize reasoning capabilities more effectively in solving character understanding tasks, aligning them more closely with the original intent behind these benchmarks.
Character Understanding
@@ 2 Related Task
Character understanding has made significant contributions to advancing story comprehension in psychological, literary, and educational research  Bower and Morrow (1990); McKee (2010); Paris and Paris (2003). This multifaceted field encompasses a variety of tasks, including character recognition, coreference resolution, summarization, role detection, goal understanding, personality analysis, and question answering Sang et al. (2022b); Bamman et al. (2020); Chen et al. (2022); Frermann et al. (2018); Vallurupalli et al. (2024); Yu et al. (2023b). Recent studies have shown promising performance by large language models (LLMs) on character understanding tasks Li et al. (2023); Stammbach et al. (2022); Goyal et al. (2023). However, these studies often overlook the potential influence of memorization on model performance. Fictional works, including movie and TV show scripts, famous novels, and news texts, serve as primary sources for narrative understanding Sang et al. (2022a); Tapaswi et al. (2016); Caselli and Vossen (2017). These texts, much like copyrighted books Chang et al. (2023a), are prone to strong memorization. In our work, we measured the verbatim capabilities of LLMs on these sources, as well as the potential inflation of benchmark results caused by overly strong verbatim memory. Our approach also mitigates this effect.
Verbatim and Gist Memory
Cognitive science defines two types of memory: verbatim memory and gist memory Brainerd and Reyna (2002); Reyna and Brainerd (1998). Verbatim memory recalls the exact wording of a sentence, such as "The black-and-white cat sat on a soft, red mat in the living room," retaining every detail about the cat’s color, the mat’s texture and color, and the location. In contrast, gist memory captures the essential meaning, like "A cat was sitting indoors," focusing on the core idea—the presence of a cat and its general location—while discarding details such as the cat’s color or the specific description of the mat. This highlights how verbatim memory retains all specific details, while gist memory abstracts the key information by ignoring less critical details. Cognitive science research suggests that humans primarily rely on gist memory for reasoning, and these two types of memory are independent, rather than hierarchical Brainerd and Reyna (1992). In recent studies on LLM memorization Carlini et al. (2023), the extraction of strings from a larger text is defined as a form of verbatim memory. Their solution primarily focuses on generating entirely new test samples that the model has never encountered Xie et al. (2024a). However, for character understanding tasks, using obscure or newly released datasets that the model hasn’t seen comes with high costs, and the quality and acquisition costs of the texts are difficult to control. Our work demonstrates that using gist memory provides a simple and effective way to alleviate this issue.
@@ 3 Data and Experiment Setup
In our experiments, we utilized two distinct datasets: (1) Benchmark Datasets: six pre-existing datasets designed to evaluate various character understanding tasks, and (2) Fictional Works Dataset: a newly constructed dataset consisting of script excerpts sampled directly from the original texts used to create the benchmark datasets. This dataset, drawn from 40 diverse sources—including TV shows, novels, and movies—was specifically designed to investigate GPT’s familiarity with the original texts underlying the benchmark datasets.
3.1 Benchmark Datasets
We evaluated the interplay between memory and reasoning across six types of character understanding tasks. The datasets used are detailed in this section, while task descriptions are provided in Section 6.1.
@@ 3.1 Benchmark Datasets
For coreference resolution , we use the latest released version of the MovieCoref dataset  [^1] https://github.com/usc-sail/mica-character-coref. For character guessing, we adopt the TVSHOWGUESS dataset  [^2] https://github.com/YisiSang/TVSHOWGUESS to conduct experiments. For personality understanding, we utilize the PERSONET dataset  [^3] https://github.com/Gorov/personet_acl23. For role detection, we use the CSI dataset  [^4] https://github.com/EdinburghNLP/csi-corpus. For open-domain question answering, we conduct experiments using the FriendsQA dataset  [^5] https://github.com/emorynlp/FriendsQA. For summarization task, we adopt the SummScreen dataset  [^6] https://github.com/mingdachen/SummScreen. We follow all the data preprocessing, development, and testing separation provided by the original datasets. Tasks and datasets details are in appendix G.
@@
We followed the original evaluation metrics used in the respective tasks. For classification tasks (Character Guessing, Question Answering and Personality Understanding), we report accuracy using exact match. For sequence labeling and coreference tasks (Coreference and Role Detection), we use F1 scores as the evaluation metric. For summarization tasks (Text Summarization), we report ROUGE scores to measure the overlap between the generated and reference summaries.
@@
TVshow Guessing Character Guessing TV Shows Exact Match
@@
3.2 Fictional Works Dataset
The existing six benchmark datasets, which are derived from 91 different TV shows, 12 different movies, and 33 classic literature books. We constructed the Fictional Works Dataset by uniformly sampling 200 segments from 40 selected sources, including 11 distinct TV series, 9 movies, and 20 novels, each source provides 5 scripts. Each segment contains an average of 312 words. The format of each segment is: TV Show Title; Scene, where the scene consists of a mixture of dialogue and narration from the script.
3.3 Extended Dataset Analysis
Task Name Task Type Source
Character Mining Character Identification TV Shows
LISCU Character Identification Book Reviews
MovieQA Visual Question Answering Movies
Liter Coreference Character Coreference Novels
HVV Detect Role Detection Movies, News
SYMON Visual Question Answering Movies
CEB Character Representation Novels
Love and Violence Summarization TV Shows
3.4 Model
We pass all passages with the same prompt through GPT4o (gpt-4o), [^7] https://platform.openai.com/docs/models/gpt-4-and-gpt-4-turbo using the OpenAI API. The cost for the tasks in s 4 and s 5 is approximately \$80, while the cost for the task in s 6 is approximately \$40. The character description used in section  5 is generated by GPT-4 (gpt-4-turbo) API. For the prompts and description generated by GPT4 itself, please refer to Appendix C. Due to the inherent characteristic of ChatGPT’s unstable output, even though we had set the temperature to 0, each generation may exhibit slight variations in content. In our experiments, there is a variation of 1% to 3% in the results. To mitigate this, we repeat each experiment three times and report the average scores.
4 Verbatim Memorization Evaluation
We address our first research question: How does GPT utilize and integrate multi-source information to identify the origin of a script passage?
We do this by considering a task that is focused on evaluating the extent to which models rely on verbatim memory to solve problems without requiring significant reasoning.
4.1 Task Definition
Name Source Prediction with Rationale
In this Source Memorization task, we assess the model’s ability to predict the origin of a text passage, investigating the extent to which it relies on memorized pretraining knowledge. The model is given a segment script and we ask it to identify the source while also providing a rationale for its prediction 9. This task includes two experimental settings: (1) with or without multiple-choice options, and (2) with or without the requirement for a rationale explanation, which made four combinations.
Rationale Analysis
We collected and analyzed the rationales generated by GPT, then asked GPT to performing clustering on all rationales and find groups of rationales with similar pattern. This analysis helps identify common elements the model relies on during source prediction, especially when character names are explicitly mentioned as part of the reasoning.
Effect of Input Length
To explore the impact of input length, we limited the length of text segments provided to the model. We compared its performance in two settings: one with character names and specific mentions provided, and the other with them masked. We hypothesize that verbatim memory plays a lesser role in such tasks, aligning with the human prediction process Brainerd and Reyna (1992).
4.2 Result
Source Prediction
The results presented in Table 3 reveal several notable trends across the different experimental setups. First, the model demonstrates a strong memorization ability when presented with the original script, achieving a consistent 96% accuracy even without any options and under the Chain-of-Thought (CoT) condition. This suggests a significant reliance on memorized content, as the model can accurately retrieve the source purely based on internal knowledge. Furthermore, the use of CoT reasoning substantially enhances the model’s performance across all conditions, improving the retrieval accuracy from 96% to 100% in the case of the original script. This indicates that CoT prompts help guide the model toward more accurate predictions by encouraging a more structured reasoning process. However, when character names are masked (using placeholders like P0, P1), the model’s accuracy decreases, highlighting the importance of specific names as primary retrieval cues. The drop in performance is even more pronounced when names are replaced with generic alternatives, suggesting that the model’s retrieval capability is highly dependent on the exact memorization of character names rather than broader contextual information. This is evidenced by the reduction in accuracy from 96% with the original script to 74% with replaced names. In contrast, replacing notable mentions (such as unique terms or identifiable elements other than character names) has a minimal effect on performance, with the model maintaining a high accuracy of 96%-98%. This further underscores that the model’s retrieval strategy is heavily focused on character names, and it places less emphasis on other narrative details when determining the text’s origin. Overall, these findings highlight the model’s dependency on character-specific memorization, revealing both the strengths and vulnerabilities of its retrieval mechanisms when faced with modifications to key identifying elements.
Condition With Options CoT No Options CoT
Origin Script 96% 100% 96% 98%
Name Masked (P0, P1) 92% 98% 82% 86%
Name Replaced 84% 98% 74% 80%
Notable Mentions Replaced 94% 98% 96% 96%
Name + Mention Replaced 84% 96% 72% 80%
Table 3: We find that CoT consistently improves the model’s prediction accuracy. Character names contribute significantly to accuracy, while notable mentions have a very small impact, likely due to the correlation between their influence and their frequency of occurrence in the text.
Rationale Statistics
We analyzed the 7 rationale content generated by the model across these conditions. The following table (Table 4) shows the frequency of the key rationale elements identified in the model’s output. The most frequently mentioned rationale elements include character name, content of dialogue, and setting, indicating that these factors play a significant role in the model’s reasoning process.
Rationale Element Percentage (%)
Character Name 84%
Notable Mentions 28%
Content of Dialogue 66%
Tone 40%
Humor Style 46%
Setting 46%
Social Dynamics 36%
Table 4: We found that the model primarily relies on the combinations of character names mentioned in the text, especially the main characters, when guessing the source of the text. Although notable mentions are rarely utilized, some distinctive catchphrases, special place names, and event titles are still noticed and leveraged by the model.
Word Limitation
Table 5 shows the model’s source identification accuracy with varying token limits, comparing the original script and the script with replaced names. The model consistently achieves 96% accuracy with the original script, demonstrating strong memorization even with limited context (as few as 25 tokens). In contrast, replacing character names leads to a significant drop in accuracy, especially with shorter passages. For the 25-token input, accuracy falls to 26%, highlighting the model’s heavy reliance on specific names for retrieval. Increasing the token limit to 50 tokens improves accuracy to 48%, but extending to 100 tokens shows no further gains. On average, across all 342 samples, the accuracy for replaced names recovers to 74%, suggesting some benefit from longer contexts but still falling short of the original script’s performance. This indicates that character names play a crucial role in the model’s retrieval process, with limited capacity to compensate through broader context when explicit names are obscured.
Word Limit Original Script Replaced Names
25 Words 96% 26%
50 Words 96% 48%
100 Words 96% 52%
All (Avg. 342) 96% 74%
Table 5: By leveraging verbatim memory, GPT demonstrates a surprisingly high memorization ability even with as few as 25 words. The slower improvement as the number of words increases suggests that source prediction predominantly relies on verbatim memory.
4.3 Analysis
Using text excerpts to prompt the model to guess their source highlights the model’s reliance on verbatim memory. Experiments with word count limitations revealed that the number of words had little impact on performance, further confirming that the model relies on verbatim memory for this task. By employing CoT, we encouraged the model to use gist memory for reasoning, resulting in improved performance and greater robustness against perturbations. This, in turn, provides additional evidence that the model predominantly relies on verbatim memory to solve such tasks.
5 Gist Memory Evaluation
The experiment in Section s 4 demonstrated that applying Chain-of-Thought (CoT) can prompt LLMs to utilize gist memory. In this section, we aim to explore additional methods to encourage LLMs to leverage gist memory effectively.We hypothesize that by reducing the confidence LLMs place in their verbatim memory, they are more likely to rely on gist memory while simultaneously applying reasoning. Building on previous findings that changing character names can significantly degrade performance, we intuitively believe that altering names allows the model to recognize changes in the information embedded within names, such as nationality and gender. Additionally, we pose our second question: Does name replacement affect LLMs’ reliance on gist memory, and if so, to what extent? Moreover, does the choice of name replacement method make a difference?
5.1 Task Definition
We adopt the character guessing  Sang et al. (2022b) format, and we masked the character names mentioned in the script as well. For the prompts please refer to Appendix D.

### v4

@@ Abstract
Recently, Large Language Models (LLMs) have shown impressive performance in various character understanding tasks. However, the extensive pre-training corpora used by LLMs raise concerns that they may rely on memorizing popular fictional works rather than genuinely understanding and reasoning about them. Prior work has addressed memorization in LLM reasoning tasks, but these approaches do not directly apply to character understanding. To address this, we leverage the cognitive science concepts of "gist memory" and "verbatim memory", where verbatim memory corresponds to exact recall, while gist memory captures underlying meaning. We design a hard-soft intervention framework, where the hard setting disrupts memorization by perturbing key entities, while the soft setting maintains the original text but guides the model toward reasoning. Our experiments across six character understanding tasks with GPT-4o, LLaMA3.3-70B, and DeepSeek V3 demonstrate that reducing verbatim memory reliance shifts model behavior toward reasoning. Our method leads to a performance drop of up to 45.8% when models can no longer depend on memorization, highlighting the extent to which existing benchmarks inflate performance through data contamination. Additionally, we show that prompting techniques can systematically alter a model’s response strategy, making it more dependent on either memorization or reasoning.
@@ 1 Introduction
Motivated by the concept of gist memory, we propose a prompt-based approach to mitigate model reliance on verbatim memorization in character understanding tasks. We evaluate our method on GPT-4o, LLaMA3-70B, and DeepSeek V3, introducing a hard-soft setting framework: the hard setting blocks memorization by replacing character names, while the soft setting retains the original text but employs gist-based prompts to encourage reasoning over recall. Our ablation study confirms the effectiveness of both interventions.
Building on these findings, we extend our approach to five character understanding tasks, designing task-specific gist prompts and evaluating models under both settings. Our results indicate that the Gist+NR approach best reduces memorization dependency.
In summary, we outline our study and key contributions as follows:
  1. 1.
    Character understanding tasks heavily rely on verbatim memorization. Across six benchmark tasks, all evaluated LLMs show significant performance drops when explicit memorization is disrupted, confirming that these tasks often assess recall rather than true reasoning (Table 6.2).
  2. 2.
    Gist-based prompting effectively reduces memorization dependency. Introducing gist-based prompts systematically lowers reliance on verbatim recall across all tasks, encouraging models to use reasoning instead of mechanical memorization (Section 6).
  3. 3.
    Verbatim and gist-based prompting modulate model response strategies. Models adjust their response strategies based on the prompting style—verbatim-based prompts reinforce reliance on memorization, while gist-based prompts encourage reasoning-driven answers (Section 4.3).
2 Related Work
2.1 Character Understanding and its evolution with LLM
@@
Character understanding has made significant contributions to advancing story comprehension in psychological, literary, and educational research  Bower and Morrow (1990); McKee (2010); Paris and Paris (2003). This multifaceted field encompasses a variety of tasks, including character recognition, coreference resolution, summarization, role detection, goal understanding, personality analysis, and question answering Sang et al. (2022b); Bamman et al. (2020); Chen et al. (2022); Frermann et al. (2018); Vallurupalli et al. (2024); Yu et al. (2023b). Recent studies have shown promising performance by LLMs on character understanding tasks Li et al. (2023); Stammbach et al. (2022); Goyal et al. (2023). However, these studies often overlook the potential influence of memorization on model performance. Fictional works, including movie and TV show scripts, famous novels, and news texts, serve as primary sources for narrative understanding Sang et al. (2022a); Tapaswi et al. (2016); Caselli and Vossen (2017). These texts, much like copyrighted books Chang et al. (2023a), are prone to strong memorization. Naturally, this raises curiosity about how memorization affects downstream tasks, particularly character understanding. While previous research has demonstrated the memorization capabilities of LLMs, our work further explores its specific impact on character understanding and proposes an approach to mitigate this effect.
2.2 LLMs’ Memorization and Reasoning
Among LLMs’ various capabilities Li et al. (2024c, e); Zhang et al. (2024); Li et al. (2024d), the reasoning abilities Tong et al. (2023); Yang et al. (2023); Xiong et al. (2024b); Tong et al. (2024); Xiong et al. (2024a) of LLMs Li et al. (2024b); Tan et al. (2024) have garnered significant attention, but their memorization has also been shown to play a crucial role in influencing benchmark reasoning performance Zhou et al. (2023); Xie et al. (2024a); Chen et al. (2024); Li et al. (2024a); Biderman et al. (2024). The influence of memorization is complex: data contamination can inflate performance metrics Khandelwal et al. (2020); Blevins and Zettlemoyer (2022); Carlini et al. (2023), while knowledge conflict can reduce them Su et al. (2024); Wang et al. (2024); Xie et al. (2024b); Xu et al. (2024).
While prior work has focused on detecting and mitigating memorization effects in reasoning tasks such as numerical, commonsense, and logical reasoning Xie et al. (2024a); Sainz et al. (2023); Li et al. (2024f), character understanding tasks introduce additional challenges due to their diverse formats and varying memory demands. Unlike structured reasoning tasks where excessive memorization is clearly undesirable, character understanding often requires a nuanced balance between recall and inference.
2.3 Verbatim and Gist Memory in LLM Reasoning
Cognitive science distinguishes between verbatim memory, which retains exact details, and gist memory, which captures the essential meaning Brainerd and Reyna (2002); Reyna and Brainerd (1998). For example, verbatim memory would store a sentence like "The black-and-white cat sat on a soft, red mat in the living room" with all details intact, while gist memory would reduce it to "A cat was sitting indoors," focusing on the core idea while discarding specifics. Human reasoning primarily relies on gist memory, as the two types of memory function independently rather than hierarchically Brainerd and Reyna (1992).
In LLM research, memorization is often equated with verbatim memory, where models retrieve and reproduce exact strings from pretraining data Carlini et al. (2023). A common mitigation strategy involves curating test samples that the model has never seen before Xie et al. (2024a). However, in character understanding tasks, acquiring obscure or newly released datasets can be costly and difficult to control in terms of quality. Instead, our work proposes leveraging gist memory to reduce reliance on precise memorization, encouraging LLMs to generalize character understanding beyond string-matching. By shifting the focus toward gist memory, we aim to align LLM character analysis with human-like reasoning rather than dataset recall.
@@ 3 Data and Experiment Setup
In our work, we utilize two distinct datasets to evaluate LLMs on: (1) their reasoning capabilities in character understanding under gist memory and (2) their memorization of original source texts.
3.1 Benchmark Datasets for Character Understanding Tasks
To evaluate the interplay between memorization and reasoning in character understanding tasks, we employ six publicly available datasets covering diverse task types. These datasets span different narrative mediums, including TV shows, movies, and novels, ensuring diversity in both content and character representation. The task formulations and dataset compositions are detailed below, with an overview provided in Table 1. Further task details are available in  6.1.
@@ 3.1 Benchmark Datasets for Character Understanding Tasks
We evaluate character understanding using six publicly available datasets spanning diverse sources, including TV shows, movies, and novels. For coreference resolution, we use the latest release of the MovieCoref dataset  [^1] https://github.com/usc-sail/mica-character-coref, which provides annotated coreference chains in movie scripts. For character guess, we adopt TVSHOWGUESS  [^2] https://github.com/YisiSang/TVSHOWGUESS, where models must identify the speaker of a given dialogue. For personality understanding, we use PERSONET  [^3] https://github.com/Gorov/personet_acl23, a dataset derived from literary sources with personality trait annotations. For role detection, we employ CSI  [^4] https://github.com/EdinburghNLP/csi-corpus, which focuses on identifying key characters in crime narratives. For open-domain question answering, we evaluate models on FriendsQA  [^5] https://github.com/emorynlp/FriendsQA, where questions are answered by extracting spans from TV show dialogues. Lastly, for summarization, we use SummScreen  [^6] https://github.com/mingdachen/SummScreen, which pairs TV show transcripts with corresponding summaries. We follow the original dataset authors’ preprocessing, development, and test splits to ensure comparability. Further dataset details are provided in Appendix G.
@@
We followed the original evaluation metrics used in the respective tasks, shown in Table 1.
@@
TVShow Guess Character Guessing TV Shows Exact Match
@@
3.2 Fictional Works Memorization Benchmark Dataset
This dataset is designed to evaluate LLMs’ memorization of fictional scripts by assessing their ability to recall verbatim details from diverse narrative sources.
Benchmark Sources
The six benchmark datasets described in Section 6.1 are derived from a diverse set of sources, including 91 TV shows, 12 movies, and 33 classic literary works.
Dataset Construction
To ensure broad coverage across different narrative mediums, we sample 200 script segments from 40 selected sources, comprising 11 TV series, 9 movies, and 20 novels. Each source contributes 5 excerpts, with an average segment length of 312 words.
Segment Format
Each excerpt follows the format: TV Show Title; Scene, where the scene consists of a mixture of dialogue and narration extracted from the original script.
3.3 Models
we evaluate both open-source and closed-source models, leveraging their APIs for inference. Our evaluation includes three frontier LLMs: the open-source Llama 3.3-70B AI (2024) and DeepSeek V3 Liu et al. (2024), as well as the closed-source GPT-4o OpenAI (2023).
4 Soft Setting: Memorization-Controlled Prompting in Character Guessing
First, we investigate whether prompting strategies can subtly influence LLMs’ response patterns in a soft setting, where the model is not explicitly restricted from using memorization but is instead encouraged to rely more on reasoning over direct retrieval.
4.1 Character Guessing Task Format
Let a dialogue segment be represented as $D=\{U_{1},U_{2},...,U_{n}\}$, where each utterance $U_{i}$ is paired with a speaker label $S_{i}$. In the Character Guessing Task, the model is given a dialogue excerpt without speaker annotations and must predict the correct speaker for each utterance, $f_{\text{guess}}(U_{i})\rightarrow S_{i}$. We follow the original format and setting Sang et al. (2022b), using the TVShow Guess Dataset described in Section 3.1. This task is designed to prompt the model to make guesses based on character relationships, key events, and personality traits within the text. However, its format closely resembles the name cloze tests Chang et al. (2023b) recently used to evaluate LLM memorization.
4.2 Prompting Strategies
We design two distinct prompting conditions to examine how LLMs rely on different memory mechanisms when identifying characters. Verbatim prompting encourages mechanical memorization, where the model recalls exact substrings from its training data and reconstructs missing parts without reasoning. In contrast, gist prompting focuses on extracting key information and performing memory-based matching. For character understanding tasks, the most relevant gist cues typically include character relationships, significant events, and personality traits.
All prompts share the same base instruction: "The following is a dialogue. Please identify who they are. …"
The two variations are as follows:
Verbatim Prompt: "…Do this by direct memory retrieval—do not infer based on traits, interactions, or context. Recall names exactly from past dialogue knowledge without explanations or reasoning."
Gist Prompt: "…Do this by analyzing relationships, key events, and personality traits to match them with known characters. Do not directly retrieve from memory who said the sentences."
4.3 Result and Analysis
Figure 2 visualizes the performance of the three models on the Character Guessing task under different prompting conditions. The performance trends remain consistent across models, and our key observations are as follows:
  • Baseline Performance: The baseline condition reflects the overall performance of the models, with GPT-4o demonstrating a clear lead, while Llama 3.3-70B and DeepSeek V3 exhibit similar performance.
  • Memorization Reliance: All three models show that the Verbatim condition performs nearly identically to the baseline, supporting our hypothesis that models primarily rely on memorization when answering the Character Guessing task.
  • Effect of Gist Prompting: The Gist condition results in a non-trivial performance drop, suggesting that our prompt effectively reduces the models’ reliance on memorization when generating responses.
Figure 2: The bar chart compares the performance of three LLMs under three prompting settings. Across all three models, the performance of the Verbatim and Baseline conditions is highly similar, confirming our hypothesis that the models rely on memorization when answering questions. The Gist condition, however, shows a non-trivial decline in performance, indicating that the models struggle when they cannot fully rely on memorization to generate responses.
5 Hard Setting: Memorization-Controlled Prompting under Name Replacement
Furthermore, we aim to validate whether our prompting strategies can genuinely alter the model’s reasoning patterns under a hard setting, where memorization is explicitly restricted. Previous research has suggested that name replacement may impact model performance Yu et al. (2022). In this section, we confirm this hypothesis by demonstrating that name replacement significantly affects the model’s memorization ability and quantifying its impact. Then, we examine how prompting strategies influence LLMs’ reliance on memorization versus reasoning when memorization is constrained by name replacement.
5.1 Name Replacement Strategies and Their Impact on Memorization